The Corporation purchases from suppliers in a quarter are equal to 55 percent of the next quarter's forecast sales. The payables period is 90 days. Wages, taxes and other expenses are 25 percent of sales and interest and dividends are $82 per quarter. No capital expenditures are planned.

Here are the projected quarterly sales:

Q1

Q2

Q3

Q4

Sales

1390

1520

1640

1160

Sales for the first quarter of the following year are projected at $1,740. Calculate the company's cash outlays by completing the following:

Q1

Q2

Q3

Q4

Payment of accounts

Wages, taxes and other expenses

Long-term financing expenses

(interest and dividends)

Total

Shows all the step and formula. Don't round off until you get the answer.
Answer #1

Calculate the payment outlay as follows:

Q1 $836.00 $902.00 $638.00 $957.00 $347.50 $380.00 $410.00 $290.00 Q4 Q2 Qз 2 Payment of accounts 3 Wages taxes and other exp

Formulas:

E Q3 =1160*55% =1640*25% Q2 =1520*55% 3 Wages taxes and other expenses =1390*25% Q4 =1740*55% =1160*25% Q1 =1640*55% 2 Paymen